# Secure Information Embedding and Extraction in Forensic 3D Fingerprinting

### v1

@@ 6.2 Information Embedding and Extraction
Figure 7: The reading of a fragment (cropped due to space limit), in which every bright lines is the center of a layer. The pixel values (in gray scale) are accumulated and normalized for every column and displayed in red. The peaks are marked by yellow dots, which reveal the distance between every two consecutive bright lines.
Due to the limitation of field of view, the microscope is unable to capture the entire fragment all at once. To address this, the fragment to be examined is fixed on a motorized rail slider (see Figure 5(b)). During the information extraction process, the fragment slides over the field of view of the microscope, which takes series of pictures in the meantime. Note that the pictures are taken so that every two consecutive pictures overlap. This feature allows us to fuse them together and obtain a picture of the entire fragment using ImageJ with the stitching plugin [39].
In the resulting stitched image, shown in Figure 7, the bright lines are the center of each layer. To automatically read the bits from this image, we implemented a program that reads the distance between every two adjacent lines, and places them in an array. Recall that there are two types of layers, the thin ($0.12$ mm) and the thick ($0.18$ mm). Hence, the elements in the array can be placed in three categories, i.e., two adjacent thin layers (thin-thin), a thin layer followed by a thick one or vice versa (thin-thick), and two adjacent thick layers (thick-thick). The program then applies $K$-means clustering with $K=3$ in order to differentiate the three categories.
Next, the program examine the clustering results and corrects potential errors resulting from the lighting environment during microscope reading or the image distortion in stitching. The embedding method (Figure 5) imposes certain patterns, which allow such correction. For example, a thin-thin distance must appear adjacent to another thin-thin distance, and in between thin-thin and thick-thick there must be a thin-thick.
@@
7 Evaluation
The evaluation of our prototype focuses on the information density (number of bits per unit distance, which is directly affected by the codeword length $n$ of BRC) and the recoverability of data from broken objects. The former measures the amount of information that can be embedded in an given object, and the latter provides the secure information extraction feature of SIDE.
On the theoretical side, we analyzed the growth of the codeword length $n$ as a function of $k$ and $t$, for $k\in\{64,96,128,160\}$ and $t\in\{1,2,\ldots,9\}$ (Figure 8). On the engineering side, we conduct experiments on two groups of samples, which differ in the number of embedded bits (reflected by $k$) and the number of breaks that can be tolerated (reflected by $t$).
For the first group, we printed five sample cuboids of dimension $6$mm×$\times 20$mm×$\times 90$mm, and implemented a $1$-break code with $k=98$ and $n=243$, which maps an information word of $98$ bits to a $243$ bit codeword. For each cuboid, we prepared a distinct information string and encoded it to a codeword of the aforementioned $1$-break code; the codeword is embedded in the cuboid and is decodable even if broken at most once at any arbitrary location. For the second group, we develop a $2$-break code with $k=120$ and $n=353$. The sample cuboids in this group has a dimension of $8$mm×$\times 20$mm×$\times 130$mm. The cuboids in this group are resilient to two breaks in any arbitrary location. Table I details the parameters of both groups.
For every cuboid, we broke it apart by hand into $t+1$ parts, where $t$ is the number of breaks to which it is resilient. We examined the fragments and read bits (see Section 6.2). We assumed that the orientation of the fragments (i.e., the correct $z$ direction) is known to the decoder; we postulate that closer examination of individual layers can alleviate this assumption due to the impact of gravity on the molten thermoplastic, and leave this to future work. The bit strings from the fragments were fed into the decoding algorithm, and the decoding results are compared against the original encoded data. The information extraction succeeded in all ten cases.
@@ 8 Conclusion
In this paper, we present SIDE, which aims to provide secure information embedding and extraction in forensic 3D fingerprinting against the malicious tempering of both 3D printers and the printed objects. SIDE is built upon hardware isolation mechanisms and coding-theoretic techniques. we develop break-resilient code, which allows for information extraction from the printed object even if it is broken into pieces. We leverage trusted execution environment to protect the integrity of the printing process, which guarantees the embedding of the break-resilient codeword. We implement a prototype of SIDE and validated our design.
Finally, we propose future studies in order to advance secure 3D fingerprinting technology further. On the engineering side, we propose to study the integration of trusted execution environments in commodity printers, and the trusted procurement of bits (e.g., timestamp or geolocation) for the sake of embedding. Another engineering challenge is coming up with alternative secure methods for bit embedding in order to increase information density. On the theoretical side we propose to devise practical coding methods which operate closer to the information-theoretic limits of this model [26], as well as more sophisticated methods for reconstructing the information from only a subset of fragments, rather than all of them.
9 Acknowledgment

### v3

@@ 6.2 Secure Information Extraction
Due to the limitation of field of view, the microscope is unable to capture the entire fragment all at once. To address this issue, the fragment to be examined is fixed on a motorized rail slider (see Figure 10(b)). During the information extraction process, the fragment slides over the field of view of the microscope, which takes series of pictures in the meantime. Note that the pictures are taken so that every two consecutive pictures overlap. This feature allows us to fuse them together and obtain a picture of the entire fragment using ImageJ with the stitching plugin [Stitching].
In the resulting stitched image, shown in Figure 11, the bright lines are the center of each layer. To automatically read the bits from this image, we implemented info_parser, a program that reads the distance between every two adjacent lines, and places them in an array. Recall that there are two types of layers, the thin ($0.12$ mm) and the thick ($0.18$ mm). Hence, the elements in the array can be placed in three categories, i.e., two adjacent thin layers (thin-thin), a thin layer followed by a thick one or vice versa (thin-thick), and two adjacent thick layers (thick-thick). info_parser then applies $K$-means clustering with $K=3$ in order to categorize the elements.
Next, info_parser examines the clustering results and corrects potential errors resulting from the lighting environment during microscope reading or the image distortion in stitching. The embedding method (Figure 8) imposes certain patterns, which allow such correction. For example, a thin-thin distance must appear adjacent to another thin-thin distance, and in between thin-thin and thick-thick there must be a thin-thick.
@@
Finally, the bit strings that were recovered from the fragments go through the BRC decoding procedure, whose pseudo-code is given in Algorithm 2 and implementation is made in Python with reedsolomon codec [reedsolomon], and the original information $\mathbf{w}$ is recovered.
Figure 11: The reading of a fragment (cropped due to space limit), in which every bright line is the center of a layer. The pixel values (in gray scale) are accumulated and normalized for every column and displayed in red. The peaks are marked by yellow dots, which reveal the distance between every two consecutive bright lines.
7 Security Analysis
We provide analysis on the security of SIDE framework with respect to several potential attacks. We emphasize that the goal of the adversary is not to fail the printing process. Instead, the adversary wishes to jeopardize the flow of information embedding and extraction, without affecting the printing process of the desired crime tool.
7.1 Man-in-the-Middle Attack
In the scenario where the communication between the printer control board and the printer hardware is not secure, the adversary may insert themself as a man in the middle. Specifically, the adversary hijacks the stepper events sent to the printer hardware that associate to the stepper motor which moves the nozzle along the $z$ direction. The adversary modifies the stepper events and hence strips off the embedded bits. However, since this process moves layers to incorrect heights, it results in a discrepancy between the printed tool and the 3D model in their geometric shapes, and the tool may lose its functionality.
7.2 Spoofing Attack
The adversary may conduct a spoofing attack by detaching the printer hardware from the designated control board, and operating the hardware with an alternative one which does not embeds bits in the printed tool. To tackle this attack (as well as the MIIM attack), the printer hardware must not print unless the control board identifies itself. Numerous methods have been developed to achieve this goal, e.g., cryptographic authentication. However, employing them is outside of the scope of this paper.
7.3 Denial-of-Service Attack
Since SIDE embeds information along the printing direction, the printed object must be of at least a certain height in order to hold the BRC codeword. This fact allows the adversary to conduct a denial-of-service attack on information embedding, since the 3D printer must also be available for small objects. As such, the adversary may partition the 3D models in shorter pieces and print them individually. Next, the adversary joins the pieces together with some fasteners or adhesives before committing the crime. The information embedding process is hereby bypassed.
However, we note that this process has negative impacts on the quality of the prints, since the joints are the most fragile. Indeed, the tools made by joining smaller pieces are unlikely to bear mechanically strength that is enough for criminal use. For example, a printed firearm must maintain its geometric shape after every shot under a chamber pressure, which may exceed $50,000$ psi. Thus, we do not consider the printing successful, as this attack does not persevere the functionality of the printed object.
7.4 Information-Availability Attack
In order to jeopardize information extraction, the adversary may attempt to compromise the availability of the information that can be read from the fragments. Specifically, we consider a scenario in which the adversary prints the tool, commits the crime, and damages the tool physically in order to make the embedded bits unavailable for reading. We assume that the adversary has limited time and limited access to resources, and hence unable to damage the tool in any arbitrary fashion.
Excessive Breaking: As noted in the previous section, the printed tool should be mechanically rigid enough for criminal use. As a result, with human-powered equipment such as hammer or saw, the adversary must invest intensive time efforts to excessively break the tool such that the security parameter $t$ can be exceeded. This is infeasible since the adversary has only a very limited time after committing the crime by assumption. Otherwise, the adversary may seek help from power equipment such as grinder or jigsaw, which also violates our assumption that the resources accessible by the adversary is restricted (note that household equipment, e.g., a blender, is unlikely to bear enough power to break the printed tool into fine grained pieces, even if it is printed with thermoplastic). In addition, excessive breaking (e.g., using a hammer) is likely to cause irregular breaks, i.e., which are not aligned against the printing axis. As discussed in Section 3 (see also Figure 5 and Figure 6) irregular breaks can likely be put back together using the topology of the object, leaving the de-facto number of breaks much smaller.
Surface Altering: We note that SIDE is inherently immune to this type of attacks by its design. Recall that SIDE embeds information by altering physical elements on the printing direction (e.g., layer thickness), which is an intrinsic property of the printed object and cannot be altered by post processings on the surface. Indeed, the reading of bits does not require the surface of the fragment to be free of adversarial tampering, as less-economical solutions, such as an industrial CT scanner, can be used to infer the layer thickness via tomographic analysis.
Hiding: We consider the case in which the time and accessible resource, although limited, allow the adversary to take away some small fragments from the crime scene, or equivalently, to concentrate all efforts to make bits in small intervals unreadable. For example, the adversary may pick a fragment and melt a small portion of it with a lighter. However, as long as the number of missing bits is below certain threshold, the decoding algorithm still recovers the information word $\mathbf{w}$ successfully, as shown in Corollary 1.
8 Evaluation
@@
The evaluation of our prototype focuses on the information density (number of bits per unit distance, which is directly affected by the codeword length $n$ of BRC) and the recoverability of data from broken objects. The former measures the amount of information that can be embedded in an given object, and the latter provides information extraction from broken fragments.
@@
8.1 Information Density
We first analyze the codeword length $n=|\mathbf{c}|$ provided by the BRC encoding algorithm, with respect to the security parameter $t$ and information length $k=|\mathbf{w}|$. For some $t$ and $k$, the length $n$ is the number of bits required for the codeword $\mathbf{c}$ to be break-resilient against $t$ breaks, and is inversely related to the information density. For $k\in\{64,96,128,160\}$ and $t\in\{1,2,\ldots,9\}$, the values of $n$ is shown in Figure 12.
On the engineering side, we conduct experiments on two groups of samples, which differ in the number of embedded bits (reflected by $k$) and the number of breaks that can be tolerated (reflected by $t$).
For the first group, we printed five sample cuboids of dimension $6$mm×$\times 20$mm×$\times 90$mm, and implemented a $1$-break code with $k=98$ and $n=243$, which maps an information word of $98$ bits to a $243$ bit codeword. For each cuboid, we prepared a distinct information string and encoded it to a codeword of the aforementioned $1$-break code; the codeword is embedded in the cuboid and is decodable even if broken at most once at any arbitrary location.
For the second group, we develop a $2$-break code with $k=120$ and $n=353$. The sample cuboids in this group has a dimension of $8$mm×$\times 20$mm×$\times 130$mm. The cuboids in this group is supposed to be resilient to two breaks in any arbitrary location. Table I details the parameters of both groups.
8.2 Recoverability
We continue to evaluate the recoverability or our proposed framework. For every cuboid, we broke it apart by hand into $t+1$ parts, where $t$ is the number of breaks to which it is supposedly resilient. We examined the fragments and read bits using the methods introduced in Section 6.2).
We assumed that the orientation of the fragments (i.e., the correct $z$ direction) is known to the decoder; we postulate that closer examination of individual layers can alleviate this assumption due to the impact of gravity on the molten thermoplastic, and leave this to future work. The bit strings from the fragments were fed into the decoding algorithm, and the decoding results are compared against the original encoded data. The information extraction succeeded in all ten cases.
@@ 9 Conclusion
In this paper, we present SIDE, which aims to provide secure information embedding and extraction in forensic 3D fingerprinting against the malicious tampering of both 3D printers and the printed objects, in spite of the adversarial nature of this problem.
SIDE is built upon coding-theoretic techniques and hardware isolation mechanisms. On the theoretic side, we develop break-resilient codes, which is a family of codes that allow for information extraction from the printed object even if it is broken, and even if some small pieces are lost. On the engineering side, we leverage trusted execution environment to protect the integrity of the printing process, which guarantees control-flow integrity of the embedding of the break-resilient codeword. We implement a prototype of SIDE and validated our design.
We propose future studies in order to advance secure 3D fingerprinting technology further. Although our implementation and experiment are conducted in low-cost hardware, SIDE is more suitable for high-end market since incorporating a trusted execution environment in the printer significantly increases its manufacturing cost. Hence, we propose to study the integration of trusted execution environments in commodity printers. In addition, SIDE does not include protection of the communication between the control board and the printer hardware, leaving a loophole that may be exploited to bypass secure information embedding. As such, a secure communication protocol between the two is the missing piece of the proposed framework. Moreover, the secure procurement of bits (e.g., timestamp or geolocation) for the sake of embedding from the insecure operation system is another potential research direction. On the coding-theoretical side, we propose to devise practical coding methods which operate closer to the information-theoretic limits of this model [wang2023breakresilient].
Acknowledgment
@@
[…]
Appendix A Distinct Strings
Algorithm 3 d-encode (distinct strings encoding)
1:Input: A binary string $\mathbf{w}\in{\{0,1\}}^{l\cdot m-1}$, where $l,m$ are positive integers such that $m\geq{\lceil 2\log l\rceil}+2$.
2:Output: Array dStrings of $l$ length-$m$ pairwise distinct binary strings.
3:Let $\mathbf{w}\leftarrow\mathbf{w}\circ 1$$i\leftarrow 0,i_{\text{end}}\leftarrow l-1$
4:Let $\texttt{dStrings}[a]\leftarrow\mathbf{w}[a\cdot m:(a+1)\cdot m)$ for $a\in[0,l-1]$.
5:while $i<i_{\text{end}}$ do
6:     $j\leftarrow i+1$
7:     while $j\leq i_{\text{end}}$ do
8:         if $\texttt{dStrings}[i]=\texttt{dStrings}[j]$ then
9:              $\texttt{dStrings}\leftarrow(\texttt{dStrings}[0:j-1],\texttt{dStrings}[j+1:l-1])$
10:              $j^{\prime}\leftarrow 0$
11:              repeat $j$ times:
12:                  do
13:                       $j^{\prime}\leftarrow j^{\prime}+1$
14:                  while $\exists r\in[l-2]~{}\mbox{s.t., }\texttt{dStrings}[r-1][0:{\lceil\log l\rceil}]=\textsc{binary}(j^{\prime},{\lceil\log l\rceil}+1)$
15:              $\texttt{dStrings}.\textsc{append}(\textsc{binary}(j^{\prime},{\lceil\log l\rceil}+1)\circ\textsc{binary}(i,{\lceil\log l\rceil})\circ 0^{m-2{\lceil\log l\rceil}-1})$
16:              $i_{\text{end}}\leftarrow i_{\text{end}}-1$
17:         else
18:              $j\leftarrow j+1$
19:     $i\leftarrow i+1$
20:return dStrings
Algorithm 4 d-decode (distinct strings decoding)
1:Input: Array dStrings of $l$ length-$m$ pairwise distinct binary strings.
2:Output: The information word $\mathbf{w}\in{\{0,1\}}^{k}$ used to generate dStrings.
3:while $\texttt{dStrings}[-1][-1]=0$ do
4:     $i\leftarrow\textsc{integer}(\texttt{dStrings}[l][{\lceil\log l\rceil}+2:2{\lceil\log l\rceil}+1])$
5:     $j,j^{\prime}\leftarrow\textsc{integer}(\texttt{dStrings}[l-1][1:{\lceil\log l\rceil}+1])$
6:     for all $r\in[0,l-2]$ and $s\in[j^{\prime}-1]$ do
7:         if $\texttt{dStrings}[r][1:{\lceil\log l\rceil}+1]=\textsc{binary}(s,{\lceil\log l\rceil}+1$then $j\leftarrow j-1$
8:     $\texttt{dStrings}\leftarrow(\texttt{dStrings}[0:j-1],\texttt{dStrings}[i],\texttt{dStrings}[j:l-2])$
9:$\mathbf{w}\leftarrow\texttt{dStrings}[0]\circ\cdots\circ\texttt{dStrings}[l-1]$
10:return $\mathbf{w}[:-1]$
Alg. 3 offers a procedure that maps an input word $\mathbf{w}\in{\{0,1\}}^{l\cdot m-1}$ to dStrings, which is an array of $l$ pairwise distinct binary strings of length $m$. Besides, the inverse operation, as given in Alg. 4, outputs $\mathbf{w}$ given the array dStrings.
The encoding algorithm first appends a $1$ to $\mathbf{w}$ (line 3), and as a result, $|\mathbf{w}|=l\cdot m$. The word $\mathbf{w}$ is segmented to $l$ intervals of length $m$ and placed in a tentative array dStrings (line 4). The segments are not necessarily pairwise distinct at this point, and they are referred as the original.
The encoding algorithm continues to look for identical original segments, deletes one of them, and appended a new binary string to dStrings; this new string contains the indexing information of the two identical binary strings which allows for decoding at a later point in time. Moreover, we end every new string with a 00 to distinguish it from the right most original segment which ends with a $1$.
With indices $i$ and $i_{\text{end}}$ initially set to 00 and $l-1$, the algorithm looks up every index $j\in[i+1,i_{\text{end}}]$ for a match, i.e., $\texttt{dStrings}[i]$ is identical to $\texttt{dStrings}[j]$ (line 7). Once a match is found, the latter is deleted from dStrings (line 9) and the indices $i,j$ are recorded in a binary string to be placed at the end of dStrings, which is responsible for the recovery of the deleted entry.
Note that, naïvely defining the new strings as the concatenation of binary representations of $i$ and $j$ may introduce more repeated strings; it may coincide with existing elements in dStrings. To address this issue, the algorithm starts looking for an alternative binary representation of $j$ which is not identical to the first ${\lceil\log l\rceil}+1$ bits of every existing element in dStrings.
[…]
Note that when the repetition stops, $j^{\prime}$ equals to the sum of $j$ and the number of times the condition in line 14 was true. Recall that $11\leq j\leq l-$ 1, and the latter equals to the number of unavailable slots that may be encountered during the increment of $j^{\prime}$, which is at most $l-1$ since there are only $l-1$ elements in dStrings. Therefore, $j^{\prime}\leq 2l-2$ and can be represented by ${\lceil\log l\rceil}+1$ bits.
Therefore, we use the binary representation of $j^{\prime}$ in ${\lceil\log l\rceil}+1$ to serve as the alternative representation of $j$. It is concatenated with the binary representation of $i$ (in ${\lceil\log l\rceil}$ bits since $i<l$) and sufficiently many 00’s to make a new string (line 15). The new string is appended to dStrings, and is different from every other element in the first ${\lceil\log l\rceil}+1$ bits; this fact gives the following lemma.
Lemma 4.
The new binary string being appended in line 15 is different from every existing elements in dStrings.
Lemma 4 allows us to decrease $i_{\text{end}}$ by one (line 16) since there is no need to compare $\texttt{dStrings}[j]$ with element whose index is greater than $i_{\text{end}}-1$. The algorithm terminates when there is no elements of dStrings remains to be compared (line 5), and its output satisfies the following.
Theorem 2.
Alg. 3 outputs an array dStrings of $l$ pairwise distinct binary strings of length-$m$.
Proof of Theorem 2.
Assume, for sake of contradiction, that there exist $a,b\in[0,l-1]$ and $a<b$ such that $\texttt{dStrings}[a]=\texttt{dStrings}[b]$. There are two possible cases for $\texttt{dStrings}[b]$.
If $\texttt{dStrings}[b]$ is a new string constructed in line 15, then it is distinct from every other elements on its left by Lemma 4, a contradiction. If $\texttt{dStrings}[b]$ is not a new string, then the $\texttt{dStrings}[a]$ on its left is not as well. As such, $\texttt{dStrings}[b]$ would have been deleted in line 9 when $i=a$ and $j=b$, a contradiction. ∎
We proceed to introduce the decoding procedure in Algorithm 4, which is essentially the inverse operation of the encoding process. Given the array dStrings, the decoding algorithm reads $i,j^{\prime}$ from the rightmost element if it is a constructed new string, i.e., if last bit of which is 00 (line 3). Recall that a new string is created when two identical strings is found, with $i$ being the index of the first one (the reference) and $j^{\prime}$ being the alternative index of the second (the referent).
Line 4 reads the value of $i$, and line 5 reads the value of $j^{\prime}$. Recall that $j^{\prime}$ is the index of the $j$-th available slots in a row. Hence, the variable $j$ is initially set to $j^{\prime}$, and then subtracted by the number of unavailable slots with indices less than $j^{\prime}$ to reaches the actual index of the referent (line 67). Together, $i,j$ enable the recovery of the referent, and the rightmost element is deleted (line 8). Once all new strings have been consumed in the aforementioned process, the decoding is concluded and $\mathbf{w}$ is returned (line 10).